OpenStudy (anonymous):

How many integers between 1 and 100 inclusive are either multiples of 2 or are multiples of 5?

OpenStudy (anonymous):

There are 50 multiples of 2 and 20 multiples of 5; but, some of the numbers are multiples of both. So, the answer wouldn't be 70. In that case we'd be counting integers like 10 multiple times.

OpenStudy (fibonaccichick666):

right, so what do you know about multiples of 5?

OpenStudy (fibonaccichick666):

specifically, what are the last digits of any number divisible by 5?

OpenStudy (loser66):

50 numbers from 0 to 100 divided by 2, among them, the number whose last digit is 0 divided by 5. 20 numbers divided by 5, among them 10 numbers has last digit is 0, hence take off those 10, we have 60 numbers are either divided by 2 or 5 For example, 10 divided by both 2 and 5, hence if we count it on 2, we don't count it on 5

OpenStudy (amistre64):

i had attempted to devise a prime counting function based on this type of idea

OpenStudy (amistre64):

\[\pi(n)=floor(\frac{n}{2})+floor(\frac{n}{3})-floor(\frac{n}{2*3})+...\]